It seems easy to answer but just want to know if my answer I did was right.
Question: Suppose you have a lemonade stand, and when you charge $2 per cup of lemonade you sell 120 cups. But when you raise your price to $3 you only sell 60 cups. Write an equation for the number of cups you sell as a function of the price you charge. Denote "C" for number of cups, and "P" for the price you charge. Assume the function is linear.

Answer by stanbon(75887)

Suppose you have a lemonade stand, and when you charge $2 per cup of lemonade you sell 120 cups. But when you raise your price to $3 you only sell 60 cups. Write an equation for the number of cups you sell as a function of the price you charge. Denote "C" for number of cups, and "P" for the price you charge. Assume the function is linear.
---------------------
You have two points: (2,120), (3,60)
Find the slope = [120-60]/[2-3] = -60
--------------
Find the intercept using c=60 when p=3 and m=-60:
60 = -60*3+b
60 = -180 + b
b = 240
------------
EQUATION:
C = -60p + 240
====================
Cheers,
Stan H.